find values of m and b in the following system so that the solution to the system is (-3,4)
5x+7y=b and mx+y=22

Answer by pizza(14)

Well, intuitively, one is supposed to see that the question is asking you to find the equation of two lines such that they intersect at (-3,4). This means we can actually treat each equation separately, and find b such that the line goes through (-3,4) and then find m similarly. If you have studied the equation of lines, then you should be on your way now. Otherwise...

First, 5x + 7y = b
Substitute x = -3, y = 4 into the equation
to get 5(-3) + 7(4) = 13 = b
So b=13.

Secondly, mx+y =22
Substitute x and y as before,
to get m(-3) + 4 = 22
which becomes m = 18/(-3) = -6
So m = -6